# A Greedy Quantum Route-Generation Algorithm

### v1

@@ Abstract
Routing and scheduling problems with time windows have long been important optimization problems for logistics and planning. Many classical heuristics and exact methods exist for such problems. However, there are no satisfactory methods for generating routes using quantum computing (QC), for mainly two reasons: inequality constraints, and the trade-off of feasibility and solution quality. Inequality constraints are typically handled using slack variables; and feasible solutions are found by filtering samples. These challenges are amplified in the presence of noise inherent in QC. Here, we propose a greedy algorithm that generates routes by using information from all samples obtained from the quantum computer. By noticing the relationship between qubits in our formulation as a directed acyclic graph (DAG), we designed an algorithm that adaptively constructs a feasible solution. We prove its convergence to a feasible solution, and illustrate its efficacy by solving the Fleet Sizing Vehicle Routing Problem with Time Windows (FSVRPTW). Our computational results show that this method obtains a lower objective value than the current state-of-the-art annealing approaches, both classical and hybrid, for the same amount of time using D-Wave’s Hybrid Solvers. We also show its robustness to noise on D-Wave’s Advantage 4.1 through computational results as compared to the filtering approach on DWaveSampler, even when the filtering approach is given a longer annealing time, and a larger sample size.
Index Terms:
@@
QUBO, quantum annealing, VRP, routing, greedy, D-Wave.
I Introduction
One of the most important problems in operations research is servicing customers within a certain time window. The simplest version of this is the travelling salesman problem with time windows, but more complex versions of this problem have also become relevant as transportation and logistics problems have become more complex. In this paper we study the Fleet Size Vehicle Routing Problem with Time Windows (FSVRPTW), which determines the minimum number of vehicles required to service all customers within their respective time windows.
@@ I Introduction
The problem is practically-relevant on its own, but it is also a sub-problem in many solution methods for more complex routing problems such as column generation, bender’s decomposition, branch-and-price. However, in the aforementioned solution approaches, an exact solution method to the sub-problem is not always necessary - but a computationally-tractable one, is. Therefore, the community would benefit from faster solutions to this problem.
@@
Quantum computing (QC) offers exciting possibilities to perform certain computational tasks better than classical solvers and may be able to speed up challenging optimization problems such as route planning. Certain algorithms such as Shor’s factorization or Grover’s search have proven advantage over their classical counterparts in fault-tolerant conditions. In the near future, it is not certain whether a quantum advantage can be observed due to the noise, and current size capabilities of quantum computers. As such, many hybrid approaches, involving classical and quantum computing, have been proposed. However, it still remains relevant to determine the best quantum approach to certain problems, so that when the QC capabilities become more advanced, the solution methods are readily available.
@@
Routing problems on quantum computers have been solved since Feld et al [1]. However, no annealing-based methods have shown to be practical for generating routes themselves. Existing approaches that have shown to be capable of solving problems at scale generate routes classically, and then solve a set cover or exact cover problem. On the other hand, the existing annealing-based approaches do not scale well, due to slack variables, and other inefficient encoding schemes [2]. We review them briefly: [3] solves the traveling salesperson problem with time windows (TSPTW) using slack variables and one-hot encoding. This makes the number of qubits impractical. [4] uses position-based indexing. While this works under the single-agent setting in TSPTW, it loses much in optimality in a multi-agent setting because it assumes each agent makes a maximum number of stops. [5] creatively discretizes time using a state-based description, but the formulation of time-windows has a discrepancy with the capacity and state description. [6] presents several QUBO formulations but only solves small-sized problems, and, aside from the route-based formulation has difficulty obtaining feasible solutions. [7], [8] and [9] all present creative hybrid solution methods, but still generate routes classically.
@@
The majority of existing QC solutions to Sherrington-Kirkpatrick (SK) Ising models involve sampling the QC a number of times, and eventually selecting the sample with the lowest energy. However, emerging research has shown the potential to use all of the samples from the QC, to take advantage of the correlation between samples [10]. We review them briefly here. [11] uses an iterative approach that fixes qubits based on their one-body expectation value, computed by using all of the samples returned from the annealer. Qubits with expectation values above a certain threshold are fixed to a particular value based on a majority vote, and a smaller SK Ising model is solved at the next iteration. The expectation value in some sense measures how certain we are of its value. [12], unlike [11], this paper has two separate procedures for selecting which qubits to fix, and the values they are fixed to. After selecting qubits, their values are determined by brute-force evaluation of the expectation value of the objective function, over all possible values of the selected variables. [13] uses two phases, instead of an iterative approach: 1) Run QAOA. 2) Compute the correlation matrix given by the output from sampling the QC circuit constructed from QAOA. The values of the principal eigenvector are then rounded to obtain a final solution. Other efforts to improve solution quality, e.g. unbalanced penalization [14], still cannot guarantee a feasible solution, and require heavy classical pre-processing for tuning penalty parameters. They also do not disclose computation time in their computational results.
@@
However, these approaches do not directly apply to solve constrained optimization problems; the process of fixing variables may decrease the expectation value of the energy, but it might also violate constraints. The fixed values are maintained permanently throughout the remainder of the algorithm. For constrained problems, we require a delicate approach for ensuring a feasible solution can still be returned after the variables are fixed. We propose an algorithm which takes advantage of both the emerging research to utilize the correlation between samples, and also is guaranteed to return a feasible solution to the FSVRPTW. In this paper we present an annealing-based algorithm for generating routes under time window constraints that is suitable for the NISQ era [15] by its convergence to a feasible solution, and competitive computation time and solution quality. Our main contributions are as follows:
@@
  • Present an annealing-based algorithm for generating routes, that represents the samples returned by the annealer as a DAG.
@@
  • Prove that it converges to a feasible solution.
@@
  • Show computational results on D-Wave, benchmarking it against classical, hybrid, and quantum annealing-based approaches based on computation time, and optimality, for a variety of practical-sized benchmark instances.
II Preliminaries
II-A Quadratic Unconstrained Binary Optimization
@@
Many discrete optimization problems can be formulated as a Quadratic Unconstrained Binary Optimization (QUBO) problem. It is the common format used to input optimization problems into quantum annealers. Given square matrix $Q\in\mathbb{R}^{n\times n}$, the problem is to minimize $x^{T}Qx$ over binary variables $x_{i}\in\{0,1\}^{n}$:
@@
Coefficients for linear terms in an optimization problem may be represented in the QUBO form by the diagonal elements of the matrix $Q$, since $x^{2}_{i}=x_{i}$. To represent this optimization problem on a quantum device, the QUBO is transformed into an Ising model with the Hamiltonian consisting of a weighted sum of tensor products of Pauli $Z$ operators. We use $Z_{i}$ to denote both the spin variables, and the Pauli $Z$ operators when clear from context. We may convert binary variables $x_{i}$ into spin variables $Z_{i}\in\{-1,1\}$ by $x_{i}=\frac{Z_{i}+1}{2}$ and represent them in an Ising model of the following Hamiltonian
@@
where $J_{ij},v_{i}$ encode the elements in the matrix $Q$. $C$ is a constant offset which is typically ignored since it doesn’t affect the optimal objective value. There is a one-to-one correspondence between the spin variables $Z_{i}$ and the binary variables $x_{i}$. Once a problem is cast as a QUBO, it may be solved on an annealer. A quantum annealer works by exploiting quantum mechanical effects to find low energy states of the Hamiltonian in equation (2) [16]. The output of an annealer is non-deterministic, so it is sampled many times to obtain a representative distribution. Given $M$ samples, the expected value of a variable $Z_{i}$ is $<Z_{i}>=\frac{\sum^{M}_{m}Z^{m}_{i}}{M}$, where $m$ is the index of the sample.
III Problem Formulation
We are given a timetable of customers, each with a time window $[e_{i},l_{i}]$, and a service time $q_{i}$; an example is shown in Table I. The start of service must being within the time window.
TABLE I: Timetable and distance matrix
@@ III Problem Formulation
We start by discretizing time such that for each time window, there exists a time discretization $t\in[e_{i},l_{i}]$. If an agent arrives at at customer before $e_{i}$, we allow for waiting. We create variables $x_{i,s,j,t}$ representing ”an agent leaves customer $i$ at time $s$, and subsequently leaves customer $j$ at time $t$”. We duplicate the depot to create original depot 00, and final depot $N$. The set $T$ is the set of time discretizations, the set $W$ is the set of customers, which excludes depots. For any variable $x_{i,s,j,t}$ we can compute $b_{ij}=\max\{e_{j},s+d_{ij}\}$ where $d_{ij}$ is the distance between customers $i$ and $j$. $b_{ij}$ is the earliest possible time we can start servicing customer $j$, having left customer $i$ at time $s$. We use an arc-based formulation from [6]: the constrained binary optimization (QUBO) is shown in equation (3).
The objective is in (3a). Constraint (3b) ensures that all customers are visited. Constraint (3c) enforces flow between consecutive customers. The remaining constraints are for pre-processing, and domain constraints.
$\min_{x_{i,s,j,t}}$ $\sum_{j\in W}\sum_{t}x_{0,0,j,t}$ (3a)
@@
$x_{0,s,N,t}=0,\,\forall s,t$ (3l)
$x_{i,s,j,t}\in\{0,1\}$ (3m)
III-A Pre-processing
Our formulation initially may appear to have a formidable number of variables. However, by making a few observations, we are able to prune many of the variables by fixing their values to 00, and removing them from the model. The variables we can prune correspond to constraints (3d)-(3l), and are described in Table II.
TABLE II: Constraint and Description for Pre-processing variables in problem (3)
@@ III-A Pre-processing
(3f) Can only leave original depot at time 0
@@
IV Methods
IV-A Greedy Quantum Route-Finding Algorithm
Our algorithm for solving problem (3) is shown in Algorithm 1. We first observe that any subset of the variables in our formulation can be represented as a directed acyclic graph. By combining this with the samples obtained from the annealer, we can converge quickly to a feasible solution. The main idea, is the following: the annealer recommends a set of variables to be used in constructing a DAG, and a classical routine is used to find feasible sub-paths within the DAG. The variables corresponding to customers along each of the paths are pruned. Variables that haven’t been pruned are still ”active”. At each iteration the set of active variables $X$, gets smaller, while the cumulative set of feasible paths returned $S$, gets larger. The algorithm terminates when $X^{l}=\emptyset$, although in practice may be terminated when the number of active variables is small enough to solve the problem exactly.
We use $Q$ to denote the new set of edge-and-vertex-disjoint paths found in the DAG, and $P$ to denote a single path in $Q$. Paths consist of a sequence of tuples $\{(i^{P}_{k},s_{k})\}_{k=0}^{|P|}$. We use $i^{P}_{k}$ to denote the customer at index $k$ of path $P$, and no superscript is used when clear from context. Let $X^{l}$ denote the set of variables ”active” at iteration $l$, $S^{l}$ denote the cumulative set of paths found in iterations prior to $l$, and $Q^{l}$ the set of paths found at iteration $l$. We use (3b, $i$) to denote the coverage constraints for customer $i$, and (3c, $i,s$) to denote the flow constraints for customer $i$ at timestep $s$. DAG($X$) denotes the DAG formed by the variables in set $X$. QUBO($X$) denotes solving problem (3) using annealing, over the set of variables $X$. $\mathcal{X}$ denotes the samples returned by the annealer. $A,M$ are annealing time and number of samples, resp. They are implicitly inputs to QUBO($X$). When clear from context, $i$ is the index of a single variable in $X^{l}$. We use $R^{\frown}R^{\prime}$ to denote the concatenation of two paths $R$ and $R^{\prime}$. The final set of paths returned as a solution to the FSVRPTW (3) is $S$.
Algorithm 1 Greedy Quantum Route-Finding Algorithm
1:$X^{0}=\{x_{i,s,j,t}\}$ initially, all variables ”active”, $\theta$, $M$, number of annealing samples, $A$, annealing time.

### v2

# A Greedy Annealing-based Route Generation Algorithm

@@ Abstract
Routing and scheduling problems with time windows have long been important optimization problems for logistics and planning. Many classical heuristics and exact methods exist for such problems. However, there are no satisfactory methods for generating routes using quantum computing (QC), mainly because of two reasons: inequality constraints and the trade-off of feasibility and solution quality. Inequality constraints are typically handled using slack variables, and samples are filtered to find feasible solutions. These challenges are amplified in the presence of noise inherent in QC. Here, we propose a greedy algorithm for generating routes using information from all samples obtained from the quantum computer. By noticing the relationship between qubits in our formulation as a directed acyclic graph (DAG), we designed an algorithm that adaptively constructs a feasible solution. We prove its convergence to a feasible solution (without post-processing) and an optimal solution if an exact solver solves the QUBO subproblem. We demonstrate its efficacy by solving the Fleet Sizing Vehicle Routing Problem with Time Windows (FSVRPTW). Our computational results show that this method obtains a lower objective value than the current state-of-the-art annealing approaches, both classical and hybrid quantum-classical, for the same amount of time using D-Wave’s Hybrid Solvers. We also show its robustness to noise on D-Wave’s Advantage2 through computational results as compared to the standard filtering approach on DWaveSampler.
@@
Keywords QUBO  $\cdot$ quantum annealing  $\cdot$ VRP  $\cdot$ routing  $\cdot$ greedy  $\cdot$ D-Wave.
1 Introduction
@@
One of the most important problems in operations research is servicing customers within a certain time window. The simplest version of this is the traveling salesperson problem with time windows. In this paper, we study the Fleet Size Vehicle Routing Problem with Time Windows (FSVRPTW), which determines the minimum number of vehicles required to service all customers within their respective time windows. The FSVRPTW problem is practically relevant on its own, but it is also a subproblem in many solution methods for more complex routing problems, such as column generation, Bender’s decomposition, branch-and-cut (B&C), and branch-and-price (B&P). However, in the aforementioned solution approaches, an exact solution method to the subproblem is not always necessary - but a computationally tractable one is. Therefore, the community would benefit from faster solutions to this problem.
@@
Quantum computing (QC) offers exciting possibilities for performing specific computational tasks better than classical solvers and may be able to speed up challenging optimization problems such as route planning. Specific algorithms have proven advantage over their classical counterparts in fault-tolerant conditions, such Grover’s algorithm for search over unstructured databases [8] proved to improve over the best possible classical alternative and Shor’s factorization of integers into prime factors [25], better than any known classical method. In the near future, it is not certain whether a quantum advantage can be observed due to the noise and current size capabilities of quantum computers. As such, many hybrid approaches involving classical and quantum computing have been proposed. However, it remains relevant to determine the best quantum approach to certain problems so that when the QC capabilities become more advanced, the solution methods are readily available.
@@
Routing problems on quantum computers have been solved since Feld et al. [7]. However, no annealing-based methods have shown to be practical for generating routes themselves. Existing approaches that have shown to be capable of solving problems at scale generate routes classically and solve a set cover problem. On the other hand, the existing annealing-based approaches do not scale well due to slack variables and other inefficient encoding schemes [20]. Here we review them: [24] solves the traveling salesperson problem with time windows (TSPTW) using slack variables and one-hot encoding, which makes the number of qubits impractical. [22] uses position-based indexing. While this works under the single-agent setting in TSPTW, it loses much in optimality in a multi-agent setting because it assumes each agent makes a maximum number of stops. [13] creatively discretizes time using a state-based description, but the formulation of time windows has a discrepancy with the capacity and state description. [10] presents several QUBO formulations but only solves small-sized problems and, aside from the route-based formulation, has difficulty obtaining feasible solutions. [27], [21], and [3] all present creative hybrid solution methods but still generate routes classically.
@@
The majority of existing QC solutions to Sherrington-Kirkpatrick (SK) Ising models involve sampling the QC a number of times and eventually selecting the sample with the lowest energy. However, emerging research has shown the potential to use all of the samples from the QC to take advantage of the correlation between samples [19]. We review them here. [4] uses an iterative approach that fixes qubits based on their one-body expectation value, computed by using all of the samples returned from the annealer. Qubits with expectation values above a certain threshold are fixed to a particular value based on a majority vote, and a smaller SK Ising model is solved at the next iteration. The expectation value, in some sense, measures how certain we are of its value. [5], unlike [4], has two separate procedures for selecting qubits to fix and their values. After selecting qubits, their values are determined by brute-force evaluation of the expectation value of the objective function, over all possible values of the selected variables. [6] uses two phases instead of an iterative approach: 1) Run QAOA. 2) Compute the correlation matrix given by the output from sampling the QC circuit constructed from QAOA. The values of the principal eigenvector are rounded to obtain a final solution. Other efforts to improve solution quality, e.g., unbalanced penalization [16], still cannot guarantee a feasible solution and require heavy classical pre-processing for tuning penalty parameters. They also do not disclose computation time in their computational results.
@@
However, these approaches do not directly apply to constrained optimization problems; fixing variables may decrease the expectation value of the energy, but it might also violate constraints. The fixed values are maintained permanently throughout the remainder of the algorithm. For constrained problems, we require a delicate approach to ensure a feasible solution can still be returned after the variables are fixed. We propose an algorithm that takes advantage of both emerging research to utilize the correlation between samples and is guaranteed to return a feasible solution to the FSVRPTW.
In this paper, we present an annealing-based algorithm for generating routes under time window constraints that is suitable for the NISQ era [23] by its convergence to a feasible solution, and competitive computation time and solution quality.
Our main contributions are as follows:
  1. 1.
@@
    An annealing-based algorithm for generating routes that represents the samples returned by the annealer as a DAG.
  2. 2.
@@
    Prove that it converges to a feasible solution and, if an exact solver solves the subproblem, converges to the optimal solution.
  3. 3.
@@
    Show computational results on D-Wave, benchmarking it against classical, hybrid, and quantum annealing-based approaches based on computation time and optimality for a variety of practical-sized benchmark instances.
2 Preliminaries
We now review the solutions and technologies used throughout this work.
Mixed Integer Programming (MIP) Solver: An MIP solver will solve a Mixed Integer Program exactly to obtain an optimal solution using Branch-and-Bound, which has exponential complexity in worst case [9, 18]. A generic MIP is shown in equation 1. In this work the MIP solver Gurobi 11.0.1 was used used to obtain the optimal solution to Problem (4).
$\min_{x,y}$ $\mathbf{c}^{\top}\mathbf{x}+\mathbf{d}^{\top}\mathbf{y}$ (1a)
s.t. $\mathbf{A}\mathbf{x}+\mathbf{B}\mathbf{y}\leq\mathbf{b}$ (1b)
$\mathbf{x}\in\mathbb{R}$ (1c)
$\mathbf{y}\in\mathbb{Z}$ (1d)
@@ 2 Preliminaries
Quadratic Unconstrained Binary Optimization (QUBO). Many discrete optimization problems can be formulated as a Quadratic Unconstrained Binary Optimization (QUBO) problem. It is the common format used to input optimization problems into annealers. Given square matrix $Q\in\mathbb{R}^{n\times n}$, the problem is to minimize $x^{T}Qx$ over binary variables $x_{i}\in\{0,1\}^{n}$:
@@
Coefficients for linear terms in an optimization problem may be represented in the QUBO form by the diagonal elements of the matrix $Q$ since $x^{2}_{i}=x_{i}$. To represent this optimization problem on a quantum device, the QUBO is transformed into an Ising model with the Hamiltonian consisting of a weighted sum of products of spin variables [14]. We use $Z_{i}$ to denote the spin variables. We may convert binary variables $x_{i}$ into spin variables $Z_{i}\in\{-1,1\}$ by $x_{i}=\frac{Z_{i}+1}{2}$ and represent them in an Ising model of the following Hamiltonian
@@
where $J_{ij},v_{i}$ encode the elements in the matrix $Q$. $C$ is a constant offset, which is typically ignored since it does not affect the optimal objective value. There is a one-to-one correspondence between the spin variables $Z_{i}$ and the binary variables $x_{i}$. Once a problem is cast as a QUBO, it may be solved on an annealer. A quantum annealer works by exploiting quantum mechanical effects to find low energy states of the Hamiltonian in Eq. (3) [17]. The output of an annealer is non-deterministic, so samples are used to obtain a representative distribution. Given $M$ samples, the expected value of a variable $Z_{i}$ is $\langle Z_{i}\rangle=\sum^{M}_{m}Z^{m}_{i}/M$, where $m$ is the index of the sample.
Quantum Annealer (QPU): A QPU uses adiabatic evolution to solve QUBOs by gradually evolving from an easy-to-solve initial state to the problem’s final configuration, utilizing quantum tunneling to avoid local minima [12]. We used D-Wave’s Advantage solver with num_reads=1000.
Simulated Annealing Sampler (SAS): A classical metaheuristic algorithm that mimics the physical process of annealing. It leverages the Markov Chain Monte Carlow method to explore the solution space. The probability of accepting a new solution is based on the Gibbs distribution. This probability is decreased according to an annealing schedule. We use the SimulatedAnnealingSampler (SAS) provided by D-Wave, and we use the default settings [1].
LeapHybridCQMSampler (CQM): A hybrid classical/quantum annealer that is specifically designed to solve problems with constraints. It combines several classical heuristics, such as Simulated Annealing, Tabu Search, and others, with Quantum Annealing to find a solution [11]. Its algorithmic parameters include num_reads, which is the number of samples returned by the annealer. The more samples, the greater the chance of finding a solution with low energy. We set num_reads=1000.
3 Problem Formulation
We are given a timetable of customers, each with a time window $[e_{i},l_{i}]$ and a service time $q_{i}$; an example is shown in Tab. 1. The start of service must begin within the time window.
@@ 3 Problem Formulation
[…]
[table]A table beside a figure
Figure 1: Timetable and distance matrix for $N$=2 example. Initial set of variables $X^{0}$ represented as a DAG (denoted DAG($X^{0}$)), continued from example shown in the Table, labeled with possible one-body expectation values from step 8 in Alg. 1. Variables selected in step 9 in Alg. 1, using $\theta=0.5$, are marked with solid arrows. Paths found in step 10 in Alg. 1 are highlighted in blue.
We start by discretizing time such that for each time window, there exists a time discretization in $[e_{i},l_{i}]$. If an agent arrives at a customer before $e_{i}$, we allow for waiting. We create variables $x_{i,s,j,t}$ representing “an agent leaves customer $i$ at time $s$ and subsequently leaves customer $j$ at time $t$”. We duplicate the depot to create the original depot 00 and the final depot $N$. The set $T$ is the set of time discretizations, and the set $W$ is the set of customers, which excludes depots. For any variable $x_{i,s,j,t}$ we can compute $b_{ij}=\max\{e_{j},s+d_{ij}\}$where $d_{ij}$ is the distance between customers $i$ and $j$. $b_{ij}$ is the earliest possible time we can start servicing customer $j$, having left customer $i$ at time $s$. We use an arc-based formulation from [10]: the constrained binary optimization is shown in Eq. (4).
The objective is in (4a). Constraint (4b) ensures that all customers are visited. Constraint (4c) enforces flow between consecutive customers. The remaining constraints are pre-processing and domain constraints.
$\min_{x_{i,s,j,t}\in\{0,1\}}$ $\sum_{j\in W}\sum_{t}x_{0,0,j,t}$ (4a)
@@
$x_{0,s,N,t}=0,\,\forall s,t.$ (4l)
3.1 Pre-processing
Our formulation initially may appear to have a formidable number of variables. However, by making a few observations, we are able to prune many of the variables by fixing their values to 00 and removing them from the model. The variables we can prune correspond to constraints (4d)-(4l) and are described in Table 1.
Table 1: Constraint and Description for Pre-processing variables in Problem (4)
@@ 3.1 Pre-processing
(4f) Can only leave the original depot at time 0
@@
4 Methods
4.1 Greedy Quantum Route-Generation Algorithm
Our algorithm for solving Problem (4) is shown in Alg. 1. We first observe that any subset of the variables in our formulation can be represented as a directed acyclic graph. By combining this with the samples obtained from the annealer, we can converge quickly to a feasible solution. The main idea is the following: the annealer recommends a set of variables to be used in constructing a DAG, and a classical routine is used to find feasible sub-paths within the DAG. The variables corresponding to customers along each of the paths are pruned. Variables that haven’t been pruned are still “active”. At each iteration, the set of active variables, $X$, gets smaller, while the cumulative set of feasible paths returned $S$ gets larger. The algorithm terminates when $X^{l}=\emptyset$, although, in practice, it may be terminated when the number of active variables is small enough to solve the problem exactly.
We use $Q$ to denote the new set of edge-and-vertex-disjoint paths found in the DAG and $P$ to denote a single path in $Q$. Paths consist of a sequence of tuples $\{(i^{P}_{k},s_{k})\}_{k=0}^{|P|}$. We use $i^{P}_{k}$ to denote the customer at index $k$ of path $P$, and no superscript is used when clear from the context. Let $X^{l}$ denote the set of variables “active” at iteration $l$, $S^{l}$ denote the cumulative set of paths found in iterations prior to $l$, and $Q^{l}$ the set of paths found at iteration $l$. We use (4b, $i$) to denote the coverage constraints for customer $i$, and (4c, $i,s$) to denote the flow constraints for customer $i$ at timestep $s$. DAG($X$) denotes the DAG formed by the variables in set $X$. QUBO($X$) denotes solving Problem (4) using annealing over the set of variables $X$. $\mathcal{X}$ denotes the samples returned by the annealer. $A$ and $M$ are the annealing time and number of samples, respectively. They are implicitly inputs to QUBO($X$). When clear from context, $i$ is the index of a single variable in $X^{l}$. We use $R^{\frown}R^{\prime}$ to denote the concatenation of two paths $R$ and $R^{\prime}$. The final set of paths returned as a solution to the FSVRPTW (4) is $S$.
Algorithm 1 Greedy Quantum Route-Generation Algorithm
1:$X^{0}=\{x_{i,s,j,t}\}$ initially, all variables “active”, $\theta$, $M$, number of annealing samples, $A$, annealing time.